SPECTROSCOPY WITH COMB-REFERENCED DIODE LASERS

MATTHEW CICH, GARY V. LOPEZ, PHILIP M. JOHNSON AND \underline{TREVOR J. SEARS}, Department of Chemistry, Stony Brook University, Stony Brook, New York 11794; CHRISTOPHER P. MCRAVEN, Homer L. Dodge Department of Physics and Astronomy, The University of Oklahoma, Norman, OK 72019-2061.

Extended cavity diode lasers have been stabilized by locking to components of an erbium-doped fiber laser-based frequency comb with a 250 MHz comb spacing centered at 1.5μm. We find the Allan variance of the diode laser frequency relative to the single comb component to which it is locked is of the order of a few Hz. For the system as a whole, the absolute frequency accuracy is approximately 1.5 parts in 10^{12}. In order to characterize the system more completely, we have recorded saturation dip absorption spectra of several transitions in the ν_{1} + ν_{3} combination band of acetylene near 6530 cm^{−1}. We find good agreement with published absolute frequency measurements for these transitions, which have been used as secondary frequency standards in the past. Aside from extremely precise saturation dip measurements such as these, comb-stabilized lasers should permit excellent measurements of Doppler-broadened lineshapes, both to compare with theory and for analytical applications. Progress along these lines will be reported at the meeting. Acknowledgments: T. J. Sears gratefully acknowledges support from a Brookhaven National Laboratory program development grant that enabled this work and also support for research at Brookhaven National Laboratory which was carried out under Contract No. DE-AC02-98CH10886 with the U.S. Department of Energy and supported by its Office of Basic Energy Sciences, Division of Chemical Sciences, Geosciences and Biosciences.